A state space representation of a system is below, 1 1 TO La 1 b] + (0.5 y=[0 c]a9.$+ Given that a = -14, b = -2, and c=6, what is its transfer function ? U(s) in the form of 1 com s+p, stro Please derive y
Answer #1

3-61]+ [...] a = ro 0.5 y=[ oc]o niven, a s - 1 , b =-2, C = 6, Yes? = c(SI-A) 8 + D. VS) *** [:], 8-696), c-603 D= [0] -5= 0(5+2.5) + 6(-14.+0.55) s?+25 +14 →T (5)= 35 – 84 s²+25* 14 liven, 9 (s) = a,sta% sips & Po